# Scaling Frustration Index and Corresponding Balanced State Discovery for Real Signed Graphs

### v2

@@ 4.1 Sampling Spanning Trees
We use gradient descent to approximate the frustration index and balance the signed graph in linear time. We adopt the equation from Du et al. [Du et al., 2016] that calculates the imbalance of a fully signed network. The definition of structural balance in these networks is different. According to the theory of homophily, a fully signed network is balanced if every edge is positive and the corresponding vertices have the same sign. If there is a negative edge, the vertices should have different signs. Suppose that the fully signed network is balanced based on the homophily theory. In that case, the underlying signed network (ignoring vertex signs) is also balanced because the fully signed network is a generalization of the signed network [Du et al., 2016]. The equation for computing imbalance in the fully signed network is outlined in Eq. 3
@@
where $e_{ij}$ is the sign of the edge connecting vertex $i$ to vertex $j$. $\theta_{i}$ and $\theta_{j}$ are the vertex signs (1 or -1) for vertices $i$ and $j$ respectively. The equation is a d differentiable loss function that we will attempt to minimize by treating the $\theta$ signs of the vertices as latent variables. Initially, the $\theta$ variables are relaxed to continuous random variables in the range between -1 to 1. We denote these continuous variables as $\Gamma$, which is essentially a vector with a size equal to the number of vertices in the signed graphs. The equation for optimization is then:
@@
The loss function used is outlined in Eq. 4. Next, for each gradient update iteration or round, the graphL algorithm computes the loss using Eq. 3. Note that algorithm in lines 5-6 sets $\theta_{i}$ and $\theta_{j}$ to -1 if $\Gamma_{i}$ and $\Gamma_{j}$ are negative respectively and to 1 if $\Gamma_{i}$ and $\Gamma_{j}$ are positive respectively. The algorithm computes the gradients with respect to each latent variable $\Gamma_{i}$ in $\Gamma$ vector in Eq. 4, and the gradients are:
@@
where $\Gamma_{j}$ is the neighbor of $\Gamma_{i}$, and we update the values of the elements of $\Gamma$ using $\Gamma$: $\Gamma$ $\leftarrow\Gamma-\alpha\frac{\partial L}{\partial\Gamma}$. In this way, we are directly minimizing the loss that represents the level of imbalance in the signed network. We repeat the whole process until we reach a predefined number of gradient updates $\lambda$. These latent variables in $\Gamma$ should converge to be either above 0 or below 0. Finally, we loop over every edge in the network and discretize the values of $\Gamma_{i}$ and $\Gamma_{j}$ along $e_{ij}$ to be integers 1 or -1 to be assigned back in $\theta$ vector. If $\Gamma_{i}$ and $\Gamma_{j}$ along edge $(i,j)$ have values above 0, we set $n_{i}$ and $n_{j}$ to 1. Otherwise, we set them to -1. We use Eq. 3 to approximate the frustration and increase the frustration counter by $\frac{1-e_{ij}\theta_{i}\theta_{j}}{2}$ for each edge. In addition, for each edge, if it is causing an imbalance, then we flip its sign and finally return the balanced state. Algorithm 2 summarizes the steps of the complete algorithm.
@@ 6 The graphBpp Proof of Concept
The setup, implementation, and data details are presented in the appendix in Section C. We compare the proposed method to BLP baseline [Aref et al., 2016] on SNAP [Leskovec and Krevl, 2014], Konect [Kunegis, 2013], and Amazon [He and McAuley, 2016] open-source benchmarks. For the Amazon signed graphs, we ran the scalable version of graphBpp intended for large graphs, which is Algorithm 5. For other smaller graphs, we ran Algorithm 1. Running Algorithm 1 for the Amazon signed graphs would not work because the 1000 balanced states for these graphs would not fit in memory and would crash the program. Thus, we cannot use both versions of the algorithm on all signed graphs.
Table 1: SNAP frustration for 1000 iterations of graphBpp with SEVEN different tree samplers introduced in subsection 4.1, and the baseline. BFS and Hybrid consistently perform the best. We rename the signed graphs as follows: rainFall (rain), S&P 1500 (S&P), wikiElec (wikiE), wikiR (wikiRfa), epinions (epin), and slashdot (slash), and we rename the samplers: Kruskal (Krusk), Aldous-Broder (AB), and Hybrid (Hybr).
@@ 6.1 Selecting the Spanning Tree Method
We compare the timing and frustration computation of graphBpp implementation of Algorithm 1 of SEVEN different tree sampling methods and look for the most effective and efficient sampling method for the frustration index computation. Note that graphBpp runs are non-deterministic, and we run the methods multiple times. The frustration committed and completion time is always the same for smaller graphs and within 0.1% for larger graphs. We compare the findings to the BLP baseline implementation for SNAP datasets.
Table 1 summarizes the results in terms of the approximated frustration index, and Table 2 summarizes the resulting frustration index per method. BFS-spanning trees produce balanced states of minimum edge switches, and DFS-spanning trees make trees with maximum edge switches. RDFS/Kruskal/Aldous Broder’s frustration scores are slightly better due to the randomization step. BFS discovers the optimal trees for the frustration computation, but they are repetitive without inducing the shuffling of the adjacency list when exploring a node’s neighborhood.
@@ 6.2 Impact of Number of Iterations on Frustration and Timing
Here, we evaluate the efficiency of the proposed algorithm by comparing graphBpp frustration and timing if the number of iterations increases. Figure 3 shows the change in performance for the two best tree sampling methods when the number of iterations grows. More iterations will not impact BFS sampling in smaller graphs. The frustration shows a slight improvement for the larger graphs for both methods when the number of iterations increases in Figure 3. The iteration timing for each of the implemented tree-sampling techniques for graphBpp is presented in the appendix in Section H.
Figure 3: graphBpp frustration for six benchmark datasets, two spanning three sampling approaches (BFS and RDFS-BFS), and three different iteration counts.
@@
In this experiment, we compare the baseline BLP [Aref and Neal, 2021] with graphBpp implementation with breadth-first search (BFS) spanning tree sampling in 1000 iterations in terms of the frustration index and the time it takes to approximate the frustration index for 13 benchmark graphs in Table 2. The space complexity of BLP is $O(|V|^{2})$, where $|V|$ is the number of vertices on the graph. Aref et al. state that the signed graphs with up to $100,000$ edges will be solved in 10 hours [Aref and Neal, 2021]. Since our computer can store all 1000 nearest balanced states for each of these 13 signed graphs in memory, we ran the graphBpp implementation of Algorithm 1, which is the non-scalable version of the algorithm for small and medium-sized graphs. This algorithm does not handle the case when the memory is complete (when storing the balanced states) and it is slightly faster than Algorithm 5 where it employs nearest balanced state replacement after it reaches the memory limit. All external processes are closed to prevent interference with time measurements. The measurement for both methods includes the time it takes to input the file, process it, and output the results.
@@
The last four columns of Table 2 summarize our findings on the SNAP benchmark. BLP and graphBpp computation for small graphs was fast and yielded equal indices such as highland and sampson18. Both methods retrieve correct frustration indices for the three datasets. The BLP code fails for the sparse epinions (over 100,000 vertices) and produces no results (adjacency matrix cannot fit in memory and crashes the Jupyter Notebook), where graphBpp finds 1000 nearest balanced states of the graph, the most optimal one with frustration 100,450 in under 23 minutes. BLP code on the S&P1500, wikiElec, WikiRfa, and slashdot signed graphs produces heuristic frustrations. (it crashes without outputting the time for S&P1500). On the other hand, graphBpp finds a lower frustration for S&P1500 and wikiElec. In addition, to overcome the memory restrictions of extracting and saving balanced states with their associated frequencies and frustration in the frustration cloud for graphBpp, we implement Algorithm 5 and set the $CAP$ to 75% of the total RAM size. We apply it to Amazon data in Table 3. The BLP model only worked and converged for the smallest 3 Amazon signed networks, the Core5 reviews. All Amazon ratings and graphs have several vertices $|V|$ higher than 300,000, and BLP outputs a memory error before initializing the model. The algorithm attempts to construct an adjacency matrix that does not fit into memory for any graph with more than 100,000 vertices. The serialized process takes about an extra hour for each 1 million edges, and the processing time, for a fixed $CAP$, is way less than BLP across the board with an increasing number of edges and vertices, see Figure 4. The BLP algorithm converges within 30 hours for smaller signed graphs and finds the optimal frustration index. graphBpp recovers the balanced state and associated frustration index for small graphs and in minutes for under 2 million edges. The most extensive graph we have processed is Amazon books with close to 10 million vertices and over 22 million edges, and it took 19 hours to find the nearest balanced state with frustration 3,146,316.
@@ 7 The graphL Proof of Concept
We pit graphL against graphBpp for approximating the frustration index and in order to obtain better stable states. We run both heuristics on the Konect signed graphs. We use Breadth-First Search (BFS) as the tree-sampling technique for graphBpp using Algorithm 5 since we proved BFS yields minimal edge sign switches previously. Since we might obtain different results for each run for the gradient descent-based method due to the random initialization of the $\Gamma$ vector, we run the heuristic five times and choose the minimal edge sign switches produced out of them. Table 4 summarizes the results. First, we can observe that the gradient descent-based method is much more efficient across the board because the heuristic runs in linear time, and it does not have to extract and save multiple nearest balanced states in memory because only graphBpp is capable of forming the frustration cloud. Second, the gradient descent-based approach generally produces more optimal balanced states for every signed graph than that of graphBpp except Sampson, Congress, and TwitterRef. for the same number of iterations/gradient updates. However, the former comes with a downside, which is tuning the learning rate hyperparameter and finding the proper initialization. Hence, graphBpp is advantageous in the sense that it does not need any hyperparameter tuning, and it is a non-trainable algorithm. Unlike graphBpp, graphL produces only one balanced state and cannot generate multiple nearest balanced states, which are essential for computing the consensus features proposed in [Rusnak and Tešić, 2021]. These features are used in clustering and signed network analysis [Tomasso et al., 2022b].
Table 4: Comparison of the frustration index approximation and execution time using graphBpp and graphL with 1000 iterations for both on the Konect data (except TwitterRef which is not a Konect graph).
@@ 8 Conclusion
In this paper, we present two novel algorithms, graphBpp and graphL, and demonstrate their scalability and superiority on large signed graphs with millions of edges and vertices. The detailed conclusion points are at the end of this paper.
@@ Appendix B Frustration Applications
In chemistry, the stability of fullerenes, like C60 and other spherical carbon structures, can be analyzed through the concept of the frustration index [Seif and Ahmadi, 2014]. The frustration index measures how an incoherent system responds to perturbations in large-scale signed biological networks [Iacono and Altafini, 2010]. The frustration of the network has been determining the strength of agent commitment to make a decision and win the disorder in adversarial multi-agent networks [Fontan and Altafini, 2018]. In these networks, the strength is determined by measuring the social commitment of agents, mainly when there are disorder or adversarial actions. Winning the disorder is the process of overcoming the chaos caused by frustration in the network. The frustration arises from the mix of collaborative and antagonistic interactions, leading to an unbalanced signed graph. To overcome this disorder and make significant decisions, agents need a high level of social commitment. The frustration in the field of neuroplasticity assesses the development of brain networks, as studies have shown that a person’s cognitive performance and the frustration of the brain network have a negative correlation [Saberi et al., 2021a]. Physical chemists predict the protein-protein interaction using the frustration index of the protein signed network [Zhou et al., 2022]. Saberi et al. investigated the pattern for the formation of frustrating connections in different brain regions during multiple life stages [Saberi et al., 2022].
@@ C.1 Signed Graph Benchmarks
BitcoinAlpha/BitcoinOTC are the user-user trust/distrust network of trading users from the Bitcoin Alpha/OTC platforms. TwitterReferendum captures data from Twitter concerning the 2016 Italian Referendum. Different stances between users signify a negative tie, while the same stances indicate a positive link [Lai et al., 2018]. WikiElec is the network of users from the English Wikipedia that voted for and against each other in admin elections. Slashdot is the reply network of the technology website Slashdot. The vertices are users, and the edges are replies. The edges of WikiConflict represent positive and negative conflicts between users of the English Wikipedia. WikiPolitics is an undirected signed network that contains interactions between the users of the English Wikipedia that have edited pages about politics. Each interaction, such as text editing and votes, is given a positive or negative value. Epinions is the trust and distrust network of Epinions, an online product rating site. It incorporates individual users connected by directed trust and distrust links.
Amazon Signed Graphs The Amazon ratings and reviews data [He and McAuley, 2016] provides rating information between 0 (low) and 5 (high) of the Amazon users on different products. We have transformed the graphs into 18 signed bipartite graphs. The raw Amazon data was originally in .json form and maintained the rating, the item I.D., and the user I.D. Here, the user I.D.s and the item I.D.s are the nodes, and the edges between them are constructed based on the rating value. If the rating is 5 and 4, it implies a positive edge; if the rating is 3 and 2, it gives no edge; and if the rating is 0 and 1, it provides a negative edge. Table 5 summarizes the large signed graphs stemming from the process.

### v3

@@ 4.1 Sampling Spanning Trees
We use gradient descent to approximate the frustration index and balance the signed graph linearly. We adopt the equation from Du et al. [Du et al., 2016] that calculates the imbalance of a fully signed network. The definition of structural balance in these networks is different. According to the theory of homophily, a fully signed network is balanced if every edge is positive and the corresponding vertices have the same sign. If there is a negative edge, the vertices should have different signs. Suppose that the fully signed network is balanced based on the homophily theory. In that case, the underlying signed network (ignoring vertex signs) is also balanced because the fully signed network is a generalization of the signed network [Du et al., 2016]. The equation for computing imbalance in the fully signed network is outlined in Eq. 3
@@
where $e_{ij}$ is the sign of the edge connecting vertex $i$ to vertex $j$. $\theta_{i}$ and $\theta_{j}$ are the vertex signs (1 or -1) for vertices $i$ and $j$ respectively. The equation is a d differentiable loss function that we will attempt to minimize by treating the $\theta$ signs of the vertices as latent variables. Initially, the $\theta$ variables are relaxed to continuous random variables in the range between -1 to 1. We denote these continuous variables as $\Gamma$, a vector with a size equal to the number of vertices in the signed graphs. The equation for optimization is then:
@@
The loss function used is outlined in Eq. 4. Next, for each gradient update iteration or round, the graphL algorithm computes the loss using Eq. 3. Note that algorithm in lines 5-6 sets $\theta_{i}$ and $\theta_{j}$ to -1 if $\Gamma_{i}$ and $\Gamma_{j}$ are negative respectively and to 1 if $\Gamma_{i}$ and $\Gamma_{j}$ are positive respectively. The algorithm computes the gradients concerning each latent variable $\Gamma_{i}$ in $\Gamma$ vector in Eq. 4, and the gradients are:
@@
where $\Gamma_{j}$ is the neighbor of $\Gamma_{i}$, and we update the values of the elements of $\Gamma$ using $\Gamma$: $\Gamma$ $\leftarrow\Gamma-\alpha\frac{\partial L}{\partial\Gamma}$. In this way, we are directly minimizing the loss that represents the level of imbalance in the signed network. We repeat the process until we reach a predefined number of gradient updates $\lambda$. These latent variables in $\Gamma$ should converge to be either above zero or below 0. Finally, we loop over every edge in the network and discretize the values of $\Gamma_{i}$ and $\Gamma_{j}$ along $e_{ij}$ to be integers 1 or -1 to be assigned back in $\theta$ vector. If $\Gamma_{i}$ and $\Gamma_{j}$ along edge $(i,j)$ have values above 0, we set $n_{i}$ and $n_{j}$ to 1. Otherwise, we set them to -1. We use Eq. 3 to approximate the frustration and increase the frustration counter by $\frac{1-e_{ij}\theta_{i}\theta_{j}}{2}$ for each edge. In addition, if it is causing an imbalance for each edge, we flip its sign and finally return the balanced state. Algorithm 2 summarizes the steps of the complete algorithm.
@@ 6 The graphBpp Proof of Concept
The setup, implementation, and data details are presented in the appendix in Section C. We compare the proposed method to BLP baseline [Aref et al., 2016] on SNAP [Leskovec and Krevl, 2014], Konect [Kunegis, 2013], and Amazon [He and McAuley, 2016] open-source benchmarks. For the Amazon signed graphs, we ran the scalable version of graphBpp intended for large graphs, which is Algorithm 5 in the appendix. For other smaller graphs, we ran Algorithm 1. Running Algorithm 1 for the Amazon signed graphs would not work because the 1000 balanced states for these graphs would not fit in memory and would crash the program. Thus, we cannot use both algorithm versions on all signed graphs.
@@
Table 1: SNAP frustration for 1000 iterations of graphBpp with SEVEN tree samplers introduced in subsection 4.1, and the baseline. BFS and Hybrid consistently perform the best. We rename the signed graphs as follows: rainFall (rain), S&P 1500 (S&P), wikiElec (wikiE), wikiR (wikiRfa), epinions (epin), and slashdot (slash), and we rename the samplers: Kruskal (Krusk), Aldous-Broder (AB), and Hybrid (Hybr).
@@ 6.1 Selecting the Spanning Tree Method
We compare the timing and frustration computation of graphBpp implementation of Algorithm 1 of SEVEN different tree sampling methods and look for the most effective and efficient sampling method for the frustration index computation. Note that graphBpp runs are non-deterministic, and we run the methods multiple times. The frustration committed and completion time is always the same for smaller graphs and within 0.1% for larger graphs. We compare the findings to the BLP baseline implementation for SNAP datasets. The results are summarized in Table 1 regarding the approximated frustration index. Table 2 summarizes the resulting frustration index per method. BFS-spanning trees produce balanced states of minimum edge switches, and DFS-spanning trees make trees with maximum edge switches. RDFS/Kruskal/Aldous Broder’s frustration scores are slightly better due to the randomization step. BFS discovers the optimal trees for the frustration computation, but they are repetitive without inducing the shuffling of the adjacency list when exploring a node’s neighborhood.
@@ 6.2 Impact of Number of Iterations on Frustration and Timing
Here, we evaluate the efficiency of the proposed algorithm by comparing graphBpp timing if the number of iterations increases. Figure 3 shows the change in performance for the two best tree sampling methods when the number of iterations grows. More iterations will not impact BFS sampling in smaller graphs. The frustration shows a slight improvement for the larger graphs for both methods when the number of iterations increases in Figure 3. The iteration timing for each implemented tree-sampling technique for graphBpp is presented in the appendix in Section H.
Figure 3: graphBpp frustration for six benchmark datasets, two spanning tree sampling approaches (BFS and RDFS-BFS), and three different iteration counts.
@@
In this experiment, we compare the baseline BLP [Aref and Neal, 2021] with graphBpp implementation with breadth-first search (BFS) spanning tree sampling in 1000 iterations in terms of the frustration index and the time it takes to approximate the frustration index for 13 benchmark graphs in Table 2. The space complexity of BLP is $O(|V|^{2})$, where $|V|$ is the number of vertices on the graph. Aref et al. state that the signed graphs with up to $100,000$ edges will be solved in 10 hours [Aref and Neal, 2021]. Since our computer can store all 1000 nearest balanced states for each of these 13 signed graphs in memory, we ran the graphBpp implementation of Algorithm 1, which is the non-scalable version of the algorithm for small and medium-sized graphs. This algorithm does not handle the case when the memory is complete (when storing the balanced states), and it is slightly faster than Algorithm 5 in the appendix, where it employs the nearest balanced state replacement after it reaches the memory limit. All external processes are closed to prevent interference with time measurements. The measurement for both methods includes the time it takes to input the file, process it, and output the results.
@@
The last four columns of Table 2 summarize our findings on the SNAP benchmark. BLP and graphBpp computation for small graphs was fast and yielded equal indices such as highland and sampson18. Both methods retrieve correct frustration indices for the three datasets. The BLP code fails for the sparse epinions (over 100,000 vertices) and produces no results (adjacency matrix cannot fit in memory and crashes the Jupyter Notebook), where graphBpp finds 1000 nearest balanced states of the graph, the most optimal one with frustration 100,450 in under 23 minutes. BLP code on the S&P1500, wikiElec, WikiRfa, and slashdot signed graphs produces heuristic frustrations. (it crashes without outputting the time for S&P1500). On the other hand, graphBpp finds a lower frustration for S&P1500 and wikiElec. In addition, to overcome the memory restrictions of extracting and saving balanced states with their associated frequencies and frustration in the frustration cloud for graphBpp, we implement Algorithm 5 in the appendix and set the $CAP$ to 75% of the total RAM size. We apply it to Amazon data in Table 3. The BLP model only worked and converged for the smallest 3 Amazon signed networks, the Core5 reviews. All Amazon ratings and graphs have several vertices $|V|$ higher than 300,000, and BLP outputs a memory error before initializing the model. The algorithm attempts to construct an adjacency matrix that does not fit into memory for any graph with more than 100,000 vertices. The serialized process takes about an extra hour for each 1 million edges, and the processing time, for a fixed $CAP$, is way less than BLP across the board with an increasing number of edges and vertices, see Figure 4. The BLP algorithm converges within 30 hours for smaller signed graphs and finds the optimal frustration index. graphBpp recovers the balanced state and associated frustration index for small graphs and in minutes for under 2 million edges. The most extensive graph we have processed is Amazon books with close to 10 million vertices and over 22 million edges, and it took 19 hours to find the nearest balanced state with frustration 3,146,316.
@@ 7 The graphL Proof of Concept
We pit graphL against graphBpp to approximate the frustration index and obtain better stable states. We run both heuristics on the Konect signed graphs. We use Breadth-First Search (BFS) as the tree-sampling technique for graphBpp using Algorithm 5 in the appendix since we proved BFS yields minimal edge sign switches previously. Since we might obtain different results for each run for the gradient descent-based method due to the random initialization of the $\Gamma$ vector, we run the heuristic five times and choose the minimal edge sign switches produced out of them. Table 4 summarizes the results. First, we can observe that the gradient descent-based method is much more efficient across the board because the heuristic runs in linear time, and it does not have to extract and save multiple nearest balanced states in memory because only graphBpp is capable of forming the frustration cloud. Second, the gradient descent-based approach generally produces more optimal balanced states for every signed graph than that of graphBpp except Sampson, Congress, and TwitterRef. for the same number of iterations/gradient updates. However, the former comes with a downside: tuning the learning rate hyperparameter and finding the proper initialization. Hence, graphBpp is advantageous because it does not need any hyperparameter tuning and is a non-trainable algorithm. Unlike graphBpp, graphL produces only one balanced state and cannot generate multiple nearest balanced states, which are essential for computing the consensus features proposed in [Rusnak and Tešić, 2021]. These features are used in clustering and signed network analysis [Tomasso et al., 2022b].
@@
Table 4: Comparison of the frustration index approximation and execution time using graphBpp and graph with 1000 iterations for both on the Konect data (except TwitterRef, which is not a Konect graph).
@@ 8 Conclusion
This paper presents two novel algorithms, graphBpp and graphL, and demonstrates their scalability and superiority on large signed graphs with millions of edges and vertices. The detailed conclusion points are at the end of the appendix.
@@ Appendix B Frustration Applications
In chemistry, the stability of fullerenes, like C60 and other spherical carbon structures, can be analyzed through the concept of the frustration index [Seif and Ahmadi, 2014]. The frustration index measures how an incoherent system responds to perturbations in large-scale signed biological networks [Iacono and Altafini, 2010]. The frustration of the network has been determining the strength of agent commitment to make a decision and win the disorder in adversarial multi-agent networks [Fontan and Altafini, 2018]. In these networks, the strength is determined by measuring the social commitment of agents, mainly when there are disorder or adversarial actions. Winning the disorder is overcoming the chaos caused by frustration in the network. The frustration arises from the mix of collaborative and antagonistic interactions, leading to an unbalanced signed graph. Agents need high social commitment to overcome this disorder and make significant decisions. The frustration in the field of neuroplasticity assesses the development of brain networks, as studies have shown that a person’s cognitive performance and the frustration of the brain network have a negative correlation [Saberi et al., 2021a]. Physical chemists predict the protein-protein interaction using the frustration index of the protein signed network [Zhou et al., 2022]. Saberi et al. investigated the pattern for the formation of frustrating connections in different brain regions during multiple life stages [Saberi et al., 2022].
@@ C.1 Signed Graph Benchmarks
BitcoinAlpha/BitcoinOTC are the user-user trust/distrust network of trading users from the Bitcoin Alpha/OTC platforms. TwitterReferendum captures data from Twitter concerning the 2016 Italian Referendum. Different stances between users signify a negative tie, while the same stances indicate a positive link [Lai et al., 2018]. WikiElec is the network of users from the English Wikipedia that voted for and against each other in admin elections. Slashdot is the reply network of the technology website Slashdot. The vertices are users, and the edges are replies. The edges of WikiConflict represent positive and negative conflicts between users of the English Wikipedia. WikiPolitics is an undirected signed network that contains interactions between the users of the English Wikipedia that have edited pages about politics. Each interaction, such as text editing and votes, is valued positively or negatively. Epinions is the trust and distrust network of Epinions, an online product rating site. It incorporates individual users connected by directed trust and distrust links.
Amazon Signed Graphs The Amazon ratings and reviews data [He and McAuley, 2016] provides rating information between 0 (low) and 5 (high) of the Amazon users on different products. We have transformed the graphs into 18 signed bipartite graphs. The raw Amazon data was originally in .json form and maintained the rating, the item I.D., and the user I.D. Here, the user I.D.s and the item I.D.s are the nodes, and the edges are constructed based on the rating value. If the rating is 5 and 4, it implies a positive edge; if the rating is 3 and 2, it gives no edge; and if the rating is 0 and 1, it provides a negative edge. Table 5 summarizes the large signed graphs stemming from the process.